OpenStudy (anonymous):

Write the equation of the quadratic function with roots -9 and and -3 and a vertex at (-6, -1).

OpenStudy (anonymous):

use y=a(x-h)^2 +k plug in the vertex (h,k) to get y=a(x+6)^2 -1 then take on of the roots (-9,0) or (-3,0) and plug in to find a 0=a(-3+6)^2 -1 to get that a=1/9 so the equation is y=(1/9)(x+6)^2 -1